# An Agentic System for Rare Disease Diagnosis with Traceable Reasoning

### v1

@@ 1 Introduction
Developing artificial intelligence (AI) systems for rare disease diagnosis presents several inherent challenges, (i) multidisciplinary: rare diseases often manifest with complex, heterogeneous, and multisystem symptoms, requiring diagnostic models to possess multidisciplinary medical knowledge and the ability to interpret diverse patient phenotypes [15, 16]; (ii) limited cases: the scarcity of cases for individual rare diseases limits the availability of training data, making it difficult to develop robust models and increasing the risk of overfitting and catastrophic forgetting; (iii) dynamic knowledge updates: the rare disease knowledge landscape is rapidly evolving, with approximately 260 to 280 new diseases added annually, according to the International Rare Diseases Research Consortium (IRDiRC) [17]. This dynamic nature demands AI systems that are not only updatable but also capable of integrating new knowledge efficiently; (iv) transparency and traceability: clinical deployment demands interpretability: diagnostic suggestions must be accompanied by transparent, traceable reasoning to support clinician trust and accountability.
@@ 2.2 Evaluation Settings
To validate the reliability of our automated evaluation approach, we engaged eight rare disease specialists, each with over 10 years of clinical experience, to independently verify the accuracy of the LLM-based assessments. The Pearson correlation coefficient between physician rankings and LLM rankings was 0.8689, indicating strong agreement between human expert and automated evaluations. In our analysis of 240 cases, 88% demonstrated concordant assessments between physicians and the LLM evaluation system. In 10% of cases, physicians ranked the correct diagnosis higher (i.e., assigned it a better position) than the LLM-based evaluation. This pattern suggests that physicians may consider diagnoses within the same broad disease category as sufficiently correct, potentially employing more nuanced clinical judgment that recognizes subtle indicators. Conversely, the LLM evaluation system appears to apply more conservative and stringent ranking criteria, tending to require a diagnosis that is completely and precisely correct to assign a high ranking. From a validation standpoint, these findings demonstrate that large-scale LLM-based evaluation can produce reliable and methodologically sound outcomes for automated assessment purposes, while acknowledging that human clinical expertise may capture additional diagnostic nuances. Additionally, these clinical specialists evaluated the validity of the models’ diagnostic reasoning processes to ensure clinical relevance and accuracy.
@@ 2.7 Web Application
  • Systematic clinical inquiry: The system first conducts "detailed symptom inquiry", further investigating information that helps clarify the scope of organ involvement, family genetic history, and symptom progression to narrow the diagnostic range. Users may also choose to skip this step and proceed directly to diagnosis.
@@ 4 Limitations
While DeepRare demonstrates strong performance and broad applicability, several areas offer opportunities for further enhancement. First, although the current tool architecture integrates a diverse set of specialized resources and databases, it does not yet encompass the full range of potentially valuable data sources. However, the flexible design of our agentic system and MCP-like interface readily allow for future expansion and integration of additional rare disease knowledge systems and bioinformatics tools, which may further strengthen diagnostic support.
@@ 5.1 Problem Formulation
In this paper, we focus on rare disease diagnosis, where the input of a rare disease patient’s case typically consists of two components: phenotype and genotype, denoted as =$\mathcal{I}=\{\mathcal{P},\mathcal{G}\}$. Either $\mathcal{P}$ or $\mathcal{G}$ (but not both) may be an empty set $\emptyset$, indicating the absence of the corresponding input. Specifically, the input phenotype may consist of free-text descriptions $\mathcal{T}$, structured Human Phenotype Ontology (HPO) terms $\mathcal{H}$, or both. Formally, we define, $\mathcal{P}=(\mathcal{T},\mathcal{H})$, where either $\mathcal{T}$ or $\mathcal{H}$ may be empty (i.e., $\emptyset$), indicating the absence of that input modality. The genotype input denotes the raw Variant Call Format (VCF) file generated from Whole Exome Sequencing (WES).
@@
[…]
@@ 5.2 Main Workflow
[…]
@@ 5.4 External Data Sources
  • HGMD (Human Gene Mutation Database) [77]: A collection of published germline mutations in nuclear genes associated with human inherited disease.
@@ 5.5 Clinical Evaluation Dataset Curation
A systematic data preprocessing pipeline was implemented to ensure data quality and relevance. For the MIMIC-IV-Note dataset, we applied a two-stage exclusion process: first, cases without rare disease diagnoses were filtered out (n = 318,976 excluded), where rare disease classification was determined using a LLM under prompt X guidelines. Subsequently, records with incomplete patient information were removed (n = 3,633 excluded), with information completeness defined as the ability to correctly extract HPO entities that could be successfully matched to the HPO database. This filtering process resulted in 9,185 cases. Similarly, the Xinhua Hospital dataset underwent parallel filtering using identical criteria, excluding 28,150 cases without rare disease diagnoses and 8,278 cases with incomplete information, yielding 5,820 cases.

### v2

@@ 1 Introduction
Developing artificial intelligence (AI) systems for rare disease diagnosis presents several inherent challenges, (i) multidisciplinary: rare diseases often manifest with complex, heterogeneous, and multisystem symptoms, requiring diagnostic models to possess multidisciplinary medical knowledge and the ability to interpret diverse patient phenotypes [15, 16]; (ii) limited cases: the scarcity of cases for individual rare diseases limits the availability of training data, making it difficult to develop robust models and increasing the risk of overfitting and catastrophic forgetting; (iii) dynamic knowledge updates: the rare disease knowledge landscape is rapidly evolving, with approximately 260 to 280 rare genetic diseases discovered per year, according to the International Rare Diseases Research Consortium (IRDiRC) [17]. This dynamic nature demands AI systems that are not only updatable but also capable of integrating new knowledge efficiently; (iv) transparency and traceability: clinical deployment demands interpretability: diagnostic suggestions must be accompanied by transparent, traceable reasoning to support clinician trust and accountability.
@@ 2.2 Evaluation Settings
To validate the reliability of our automated evaluation approach, we engaged eight rare disease specialists, each with over 10 years of clinical experience, to independently verify the accuracy of the LLM-based assessments. The Pearson correlation coefficient between physician rankings and LLM rankings was 0.8689, indicating strong agreement between human expert and automated evaluations. In our analysis of 240 cases, 88% demonstrated concordant assessments between physicians and the LLM evaluation system. In 10% of cases, physicians ranked the correct diagnosis higher (i.e., assigned it a better position) than the LLM-based evaluation. This pattern suggests that physicians employ more nuanced clinical judgment and may recognize subtle diagnostic indicators that the automated evaluation system overlooks, while LLMs tend to apply more conservative or stringent ranking criteria. From a validation standpoint, these findings demonstrate that large-scale LLM-based evaluation can produce reliable and methodologically sound outcomes for automated assessment purposes, while acknowledging that human clinical expertise may capture additional diagnostic nuances. Additionally, these clinical specialists evaluated the validity of the models’ diagnostic reasoning processes to ensure clinical relevance and accuracy.
@@ 2.7 Web Application
  • Systematic clinical inquiry: The system first conducts “detailed symptom inquiry”, further investigating information that helps clarify the scope of organ involvement, family genetic history, and symptom progression to narrow the diagnostic range. Users may also choose to skip this step and proceed directly to diagnosis.
@@ 4 Limitations
While DeepRare demonstrates strong performance and broad applicability, several areas offer opportunities for further enhancement. First, although our current agentic architecture integrates diverse specialized medical resources and databases, it has yet to incorporate potentially valuable data sources fully. Nevertheless, the flexibility of our agentic system, combined with its MCP-like plugin interface, enables seamless future expansion and integration of additional rare disease knowledge systems and bioinformatics tools, further enhancing diagnostic support.
@@ 5.1 Problem Formulation
In this paper, we focus on rare disease diagnosis, where the input of a rare disease patient’s case typically consists of two components: phenotype and genotype, denoted as $\mathcal{I}=\{\mathcal{P},\mathcal{G}\}$. Either $\mathcal{P}$ or $\mathcal{G}$ (but not both) may be an empty set $\emptyset$, indicating the absence of the corresponding input. Specifically, the input phenotype may consist of free-text descriptions $\mathcal{T}$, structured Human Phenotype Ontology (HPO) terms $\mathcal{H}$, or both. Formally, we define, $\mathcal{P}=(\mathcal{T},\mathcal{H})$, where either $\mathcal{T}$ or $\mathcal{H}$ may be empty (i.e., $\emptyset$), indicating the absence of that input modality. The genotype input denotes the raw Variant Call Format (VCF) file generated from Whole Exome Sequencing (WES).
@@
$\{\mathcal{D},\text{\hskip 3.0pt}\mathcal{R}\}=\mathcal{A}(\mathcal{P}),$ (1)
@@ 5.2 Main Workflow
$\hat{\mathcal{P}}=\begin{cases}\text{ }a_{\text{hpo}}(\mathcal{T}),&\text{if }\mathcal{T}\neq\emptyset.\\ \text{ }\mathcal{H},&\text{otherwise}.\end{cases}$ (2)
@@ 5.5 Clinical Evaluation Dataset Curation
As shown in Figure 8, a systematic data preprocessing pipeline was implemented to ensure data quality and relevance. For the MIMIC-IV-Note dataset, we applied a two-stage exclusion process: first, cases without rare disease diagnoses were filtered out (n = 318,976 excluded), where rare disease classification was determined using an LLM under prompt 13. Subsequently, records with incomplete patient information were removed (n = 3,633 excluded), with information completeness defined as the ability to correctly extract HPO entities that could be successfully matched to the HPO database. This filtering process resulted in 9,185 cases. Similarly, the Xinhua Hospital dataset underwent parallel filtering using identical criteria, excluding 28,150 cases without rare disease diagnoses and 8,278 cases with incomplete information, yielding 5,820 cases.